# Tiled Flash Linear Attention: More Efficient Linear RNN and xLSTM Kernels

### v1

@@ Abstract
Linear RNNs with gating recently demonstrated competitive performance compared to Transformers in language modeling. Although their linear compute scaling in sequence length offers theoretical runtime advantages over Transformers, realizing these benefits in practice requires optimized custom kernels, as Transformers rely on the highly efficient Flash Attention kernels (Dao, 2024). Leveraging the chunkwise-parallel formulation of linear RNNs, Flash Linear Attention (FLA) (Yang & Zhang, 2024) shows that linear RNN kernels are faster than Flash Attention, by parallelizing over chunks of the input sequence. However, since the chunk size of FLA is limited, many intermediate states must be materialized in GPU memory. This leads to low arithmetic intensity and causes high memory consumption and IO cost, especially for long-context pre-training. In this work, we present Tiled Flash Linear Attention (TFLA), a novel kernel algorithm for linear RNNs, that enables arbitrary large chunk sizes by introducing an additional level of sequence parallelization within each chunk. First, we apply TFLA to the xLSTM with matrix memory, the mLSTM (Beck et al., 2024). Second, we propose an mLSTM variant with sigmoid input gate and reduced computation for even faster kernel runtimes at equal language modeling performance. In our speed benchmarks, we show that our new mLSTM kernels based on TFLA outperform highly optimized Flash Attention, Linear Attention and Mamba kernels, setting a new state of the art for efficient long-context sequence modeling primitives.
@@ 1 Introduction
Besides the standard recurrent execution, linear RNNs allow for a parallel formulation, which, like Attention, calculates all outputs in parallel. The parallel formulation leverages the insight from linear Attention (Katharopoulos et al., 2020), which showed that kernelized dot-product-based attention can be reinterpreted as a linear RNN with matrix-valued states. Due to the linear nature of the recurrence, the computation can be split into a recurrent part, which computes intermediate RNN states, and a parallel part, which fully utilizes the hardware for computing the outputs in between (Sun et al., 2023; Hua et al., 2022).
Yang et al. (2024b) show that their custom FlashLinearAttention (FLA) kernels based on the chunkwise-parallel formulation of linear RNNs provide faster runtimes than FlashAttention. This is achieved by first splitting the sequence into chunks and materializing the first RNN state of each chunk in GPU memory. Subsequently, in the parallel part they employ one level of sequence parallelism and compute the outputs for each chunk in parallel. For a small chunk size and long sequences, this leads to a large amount of intermediate states to be stored and loaded from GPU memory, which increases memory consumption and memory input/output (IO) cost. Since modern GPUs see a faster increase in computation throughput than memory bandwidth, it is essential to minimize large memory IO. A simple approach would be to increase the chunk size. However, the chunk size of FLA is limited by the physical SRAM available on the GPU.
To solve this problem, we introduce TiledFlashLinearAttention (TFLA) which enables unlimited chunk sizes by introducing a second level of sequence parallelism via tiling matrix computations within each chunk. This enables fast kernels and allows us to efficiently balance memory consumption and IO vs. computation.
@@ 2.1 Recurrent Formulation
In its recurrent formulation, the mLSTM cell processes the series of input vectors $\bm{x}_{t}\in\mathbb{R}^{d}$ for time steps $t\in\{1,\dots,T\}$ mapping a state $\left(\bm{h}_{t-1},\bm{C}_{t-1},\bm{n}_{t-1},m_{t-1}\right)$ to a successor state $\left(\bm{h}_{t},\bm{C}_{t},\bm{n}_{t},m_{t}\right)$ given an input $\bm{x}_{t}$. Here, $\bm{h}_{t}\in\mathbb{R}^{d_{hv}}$ denotes the hidden state, $\bm{C}_{t}\in\mathbb{R}^{d_{qk}\times d_{hv}}$ denotes the cell state responsible for long-term memory, $\bm{n}_{t}\in\mathbb{R}^{d_{qk}}$ denotes the normalizer state, and $m_{t}\in\mathbb{R}$ denotes the max state. Together normalizer and max state control the magnitude of the exponential input gate and ensure stability (see Appendix D.1).
The recurrent mLSTM formulation is given by the following state update equations:
@@ 2.2 Chunkwise-Parallel Formulation
Given the logarithmic forget gates $\bar{\bm{\mathrm{f}}}^{(k)}=\log\sigma(\tilde{\bm{\mathrm{f}}}^{(k)})\in\mathbb{R}^{L}$ and input gates $\bar{\bm{\mathrm{i}}}^{(k)}=\log\exp(\tilde{\bm{\mathrm{i}}}^{(k)})\in\mathbb{R}^{L}$, we can compute the chunkwise gates as
@@
We refer to Appendix B.2 for more details on the chunkwise gate computation. The summed forget gates $\mathrm{g}_{k}$ contain the forget gate contribution of all forget gates within a chunk. The cumulative forget gate vectors $\bm{\mathrm{b}}_{k}$ contain the forget gate contributions from the beginning of the chunk up to the current time step within the current chunk. The cumulative input gate vectors $\bm{\mathrm{a}}_{k}$ contain the input gates for every timestep as well as the forget gate contributions from the current time step to the end of the chunk.
@@
where $\bar{\mathrm{g}}_{k}$ and $\overline{\bm{\mathrm{a}}}_{k}$ are the stabilized chunkwise gates (see Appendix B.2 for details). This recurrent part resembles the fully recurrent formulation in Section 2.1, but instead of computing the intermediate states for every timestep $t$, we compute them directly for every $L$ time steps without materializing the states in between.
@@ 3 Tiled Flash Linear Attention
FlashLinearAttention (Yang et al., 2024b) introduces a fast kernel algorithm for the chunkwise formulation for Linear Attention (cf. Section 2.2 without gates) and shows that their implementation is faster than optimized FlashAttention (Dao, 2024). This speedup is achieved by single level sequence parallelism, where the states $\bm{C}_{k}$ are first materialized in GPU memory and then the outputs $\bm{\mathrm{H}}^{(k)}$ are computed in parallel.
However, since FlashLinearAttention is limited in chunk size (typically $L=64$), we have to materialize many states, where the number of states is $N_{c}=\lceil T/L\rceil$. This leads to low arithmetic intensity and high GPU memory consumption, which poses challenges especially for long-context pre-training. To address this issue, we introduce Tiled Flash Linear Attention (TFLA) in Section 3.2, which enables arbitrary large chunk sizes.
@@ 3.1 GPU Fundamentals
GPU Hierarchy.
At the lowest level the GPU runs multiple Threads, operating on small but fast register memory in parallel. On the software side usually multiple (e.g. 32) Threads are grouped together into Warps. Again, multiple Warps are grouped into Thread blocks which together execute a kernel on a physical core, called streaming multiprocessor (SM). Warps or Threads within the same Thread block can communicate data through special on-chip shared memory (SRAM). When executing a kernel, a grid (with typically 3 dimensions) of Thread blocks that run in parallel is launched on the GPU. All Thread blocks have access to the large but slow off-chip high-bandwidth memory (HBM), which has both the largest latency and least bandwidth of all GPU memory. For efficient kernels it is important to minimize memory read and writes from and to HBM.
Specialized Compute Units.
Modern GPUs have specialized compute units – called tensor cores – that accelerate matrix multiplications on GPUs. Tensor cores have most of the GPU compute and are accessed at the warp or block level. For efficient kernels it is important to maximize tensor core utilization.
@@ 3.2 More Efficient Kernels via Two Level Sequence Parallelism
TiledFlashLinearAttention (TFLA) enables fast kernels and a trade-off between memory consumption and computational efficiency by introducing two levels of sequence parallelism (see Figure 5). The first level is the parallelisation over the chunks of the sequence, which requires to compute and materialize intermediate states $\bm{C}_{k}$ in GPU HBM. For this we use a recurrent kernel similar to previous work (Yang et al., 2024b). The second level is the parallelisation within each chunk, which is achieved by tiling the intra chunk attention matrix along the chunk dimension. This second level of parallelism enables large chunk sizes and hence reduces the memory consumption for the intermediate states as we have to store and load $N_{c}=\lceil T/L\rceil$ intermediate states in HBM on each kernel call, where $T$ is the sequence length and $L$ is the chunk size. In addition to the two levels of sequence parallelisation and the naive parallelisation over the batch and head dimension, TFLA also parallelizes over the embedding dimension, resulting in a massive parallelisation over five dimensions, which is crucial for achieving high performance on modern GPUs.
@@
In simplified form, the intra-chunk parallel forward pass of the mLSTM for a chunk $k$ can be written as three matrix multiplications, which we fuse into a single kernel:
@@
Figure 2 illustrates these matrix multiplications. In order to parallelize the matrix multiplications we introduce the block sizes $B_{Lhq}$, $B_{Lkv}$, $B_{dqk}$ and $B_{dhv}$ for the attention matrix, query, key, value and hidden state dimensions $L_{hq}$, $L_{kv}$, $d_{qk}$ and $d_{hv}$, along which we either parallelize or accumulate over by using a loop inside the kernel.
For the forward pass $\bm{\mathrm{H}}^{(k)}$ kernel we parallelize over the outer sequence dimension $L_{hq}$ with $N_{Lhq}=L_{hq}/B_{Lhq}$ programs and the outer embedding dimension $d_{hv}$ with $N_{dhv}=d_{hv}/B_{dhv}$ programs. We loop over the inner dimensions $L_{kv}$ and $d_{qk}$, which are tiled by the block sizes $B_{Lkv}$ and $B_{dqk}$ respectively.
@@
For the mLSTM we cannot simply accumulate the results of the matrix multiplications $\bm{\mathrm{H}}^{(k)}_{\text{intra}}$ along the $L_{kv}$ dimension and $\bm{\mathrm{H}}^{(k)}_{\text{inter}}$ due to the stabilization of the exponential input gate with the max state $m_{t}$. The max state tracks the maximum of the forget and input gates over time and is used to stabilize the exponential input gate similar to the safe softmax computation (Milakov & Gimelshein, 2018). Since we compute the hidden state output $\bm{\mathrm{H}}^{(k)}$ in blocks along the chunk size (i.e. time) dimension $L_{kv}$, we need to rescale during accumulation of the block results for $\bm{\mathrm{H}}^{(k)}_{\text{intra}}$ and the overall results into $\bm{\mathrm{H}}^{(k)}$ in the same way as FlashAttention (Dao, 2024). We provide details on the rescaling in Section B.2. For the backward pass there is no rescaling necessary as we store the max states in the forward pass and reuse them in the backward pass.
The pseudocode for the forward pass of TFLA for the mLSTM is listed in Algorithm 1.
@@
$\delta\bm{Q}^{(k)}$ P L P L $\left(\frac{d_{hv}}{B_{dhv}},\frac{L_{hq}}{B_{Lhq}},\dots\right)$
$\delta\bm{K}^{(k)}$ L P P L $\left(\frac{d_{hv}}{B_{dhv}},\frac{L_{hq}}{B_{Lhq}},\dots\right)$
$\delta\bm{V}^{(k)}$ L P L P $\left(\frac{d_{hv}}{B_{dhv}},\frac{L_{hq}}{B_{Lhq}},\dots\right)$
@@ 4.2 Normalization of mLSTM and Linear RNNs
Motivated by the performance of mLSTMsig, we seek to understand the differences between mLSTMsig and mLSTMexp empirically. To approach this, we draw inspiration from the concept of frequency response and transfer function analysis for control systems design, where typically the amplitude ratio or gain of output and input signals for different frequencies is considered (Ogata, 2010, Ch. 7). In our case, we analyze the transfer behavior of mLSTMsig and mLSTMexp for random inputs $\bm{q}_{t},\bm{k}_{t}$ and $\bm{v}_{t}$ and different input gate and forget gate preactivation values $\tilde{\mathrm{i}}_{t}$ and $\tilde{\mathrm{f}}_{t}$.
@@
Interestingly, almost all other gated linear RNN variants also place a normalization layer after the RNN cell (Dao & Gu, 2024; Sun et al., 2023; Qin et al., 2024b; Yang et al., 2024b). Often this is justified with improved training stability, but a more thorough discussion is missing (Lieber et al., 2024). Qin et al. (2022) analyze the effect of the norm layer after a non-gated, kernel-based linear attention (Katharopoulos et al., 2020) layer and show that this effectively prevents unbounded gradients. We also confirm that the norm layer has a significant impact on training stability and the gradient norm during training. In Section 5.1 we show that initializing the input gate bias at larger negative values, as suggested by our transfer behavior analysis in Figure 3, prevents large gradient norm variance and spikes during training.
@@
We hypothesize that at this point the normalization layer does not only have a stabilizing effect by controlling the magnitude of the layer activations through rescaling, but also actively participates in the information routing or gating mechanism of the linear RNN. For example, if the squared norm of $\bm{C}_{t}^{\top}\,\bm{q}$, which is controlled by input and forget gates through $\bm{C}_{t}^{\top}$, is smaller than the epsilon, the denominator in the $\mathrm{NORM}(\bm{x})$ layer is dominated by $\epsilon$ and the output moves towards zero (indicated by the purple area in Fig. 3). Hence, by moving through the x-y plane in Fig. 3, the gates could learn to suppress or amplify any input in the sequence.
In Section D.2 we show additional experiments on the effect of varying the normalization layer epsilons and different modifications of the normalizers for the mLSTM.

### v2

@@ Abstract
Linear RNNs with gating recently demonstrated competitive performance compared to Transformers in language modeling. Although their linear compute scaling in sequence length offers theoretical runtime advantages over Transformers, realizing these benefits in practice requires optimized custom kernels, as Transformers rely on the highly efficient Flash Attention kernels (Dao, 2024). Leveraging the chunkwise-parallel formulation of linear RNNs, Flash Linear Attention (FLA) (Yang & Zhang, 2024) shows that linear RNN kernels are faster than Flash Attention, by parallelizing over chunks of the input sequence. However, since the chunk size of FLA is limited, many intermediate states must be materialized in GPU memory. This leads to low arithmetic intensity and causes high memory consumption and IO cost, especially for long-context pre-training. In this work, we present Tiled Flash Linear Attention (TFLA), a novel kernel algorithm for linear RNNs, that enables arbitrary large chunk sizes and high arithmetic intensity by introducing an additional level of sequence parallelization within each chunk. First, we apply TFLA to the xLSTM with matrix memory, the mLSTM (Beck et al., 2024). Second, we propose an mLSTM variant with sigmoid input gate and reduced computation for even faster kernel runtimes at equal language modeling performance. In our speed benchmarks, we show that our new mLSTM kernels based on TFLA outperform highly optimized Flash Attention, Linear Attention and Mamba kernels, setting a new state of the art for efficient long-context sequence modeling primitives.
@@ 1 Introduction
Besides the standard recurrent execution, linear RNNs allow for a parallel formulation, which, like Attention, calculates all outputs in parallel. The parallel formulation leverages the insight from linear Attention (Katharopoulos et al., 2020), which showed that kernelized dot-product-based attention can be reinterpreted as a linear RNN with matrix-valued states. Due to the linear nature of the recurrence, the computation can be split into a recurrent part, which computes intermediate RNN states, and a parallel part, which fully utilizes the hardware for computing the outputs in between these intermediate states (Sun et al., 2023; Hua et al., 2022).
Yang et al. (2024b) show that their custom FlashLinearAttention (FLA) kernels based on the chunkwise-parallel formulation of linear RNNs provide faster runtimes than FlashAttention. This is achieved by first splitting the sequence into chunks and materializing the first RNN state of each chunk in GPU memory. Subsequently, in the parallel part they employ one level of sequence parallelism and compute the outputs for each chunk in parallel. For a small chunk size and long sequences, this leads to a large amount of intermediate states to be stored and loaded from GPU memory, which increases memory consumptionand decreases arithmetic intensity. Since modern GPUs see a faster increase in computation throughput than memory bandwidth (Gholami et al., 2024), it is essential to minimize large memory IO and increase arithmetic intensity. A simple approach would be to increase the chunk size. However, the chunk size of FLA is limited by the physical SRAM available on the GPU.
To solve this problem, we introduce TiledFlashLinearAttention (TFLA) which enables unlimited chunk sizes by introducing a second level of sequence parallelism via tiling of the matrix computations in sequence dimension within each chunk. This increases the arithmetic intensity of the kernels and allows us to efficiently balance memory consumption and IO vs. computation.
@@ 2.1 Recurrent Formulation
In its recurrent formulation, the mLSTM cell processes the series of input vectors $\bm{x}_{t}\in\mathbb{R}^{d}$ for time steps $t\in\{1,\dots,T\}$ mapping a state $\left(\bm{h}_{t-1},\bm{C}_{t-1},\bm{n}_{t-1},m_{t-1}\right)$ to a successor state $\left(\bm{h}_{t},\bm{C}_{t},\bm{n}_{t},m_{t}\right)$ given an input $\bm{x}_{t}$. Here, $\bm{h}_{t}\in\mathbb{R}^{d_{hv}}$ denotes the hidden state, $\bm{C}_{t}\in\mathbb{R}^{d_{qk}\times d_{hv}}$ denotes the cell state responsible for long-term memory, $\bm{n}_{t}\in\mathbb{R}^{d_{qk}}$ denotes the normalizer state, and $m_{t}\in\mathbb{R}$ denotes the max state. Together normalizer and max state control the magnitude of the exponential input gate and ensure stability (see Appendix D.1). The recurrent mLSTM formulation is given by the following state update equations:
@@ 2.2 Chunkwise-Parallel Formulation
Given the logarithmic forget gates $\bar{\bm{\mathrm{f}}}^{(k)}=\log\sigma(\tilde{\bm{\mathrm{f}}}^{(k)})\in\mathbb{R}^{L}$ and input gates $\bar{\bm{\mathrm{i}}}^{(k)}=\log\exp(\tilde{\bm{\mathrm{i}}}^{(k)})\in\mathbb{R}^{L}$, we can compute the logarithmic chunkwise gates as [^1] See Appendix B.2 for more details on the chunkwise gates.
@@
Figure 2: Illustration of the chunkwise gates $\bm{\mathrm{a}}_{k}$, $\bm{\mathrm{b}}_{k}$ and $\mathrm{g}_{k}$ with chunk size $L=4$. Each arrow denotes an element in the gate vectors. See Figure 8 in Appendix B.2 for more details.
In Figure 2, we show the summed forget gates $\mathrm{g}_{k}$ contain the forget gate contribution of all forget gates within a chunk. The cumulative forget gate vectors $\bm{\mathrm{b}}_{k}$ contain the forget gate contributions from the beginning of the chunk up to the current time step within the current chunk. The cumulative input gate vectors $\bm{\mathrm{a}}_{k}$ contain the input gates for every timestep as well as the forget gate contributions from the current time step to the end of the chunk.
@@
where $\bar{\mathrm{g}}_{k}$ and $\overline{\bm{\mathrm{a}}}_{k}$ are the stabilized chunkwise gates.11footnotemark: 1 This recurrent part resembles the fully recurrent formulation in Section 2.1, but instead of computing the intermediate states for every timestep $t$, we compute them directly for every $L$ time steps without materializing the states in between.
@@ 3 Tiled Flash Linear Attention
FlashLinearAttention (Yang et al., 2024b) introduces a fast kernel algorithm for the chunkwise formulation for Linear Attention (cf. Section 2.2 without gates) and shows that their implementation is faster than optimized FlashAttention (Dao, 2024). This speedup is achieved by single level sequence parallelism, where the states $\bm{C}_{k}$ are first materialized in GPU memory and then the outputs $\bm{\mathrm{H}}^{(k)}$ are computed in parallel. However, since in FlashLinearAttention the chunk size parameter determines the tile sizes in SRAM, the maximum chunk size is limited (typically $L=64$) by the physical SRAM size of the GPU. Therefore, we have to materialize many states in HBM, where the number of states is $N_{c}=\lceil T/L\rceil$. This leads to low arithmetic intensity and high GPU memory consumption, which poses challenges especially for long-context pre-training. To address this issue, we introduce TiledFlashLinearAttention (TFLA) in Section 3.2, which enables arbitrary large chunk sizes.
@@ 3.1 GPU Fundamentals
GPU Hierarchy. At the lowest level the GPU runs multiple Threads, operating on small but fast register memory in parallel. On the software side usually multiple (e.g. 32) Threads are grouped together into Warps. Again, multiple Warps are grouped into Thread blocks which together execute a kernel on a physical core, called streaming multiprocessor (SM). Warps or Threads within the same Thread block can communicate data through special on-chip shared memory (SRAM). When executing a kernel, a grid (with typically 3 dimensions) of Thread blocks that run in parallel is launched on the GPU. All Thread blocks have access to the large but slow off-chip high-bandwidth memory (HBM), which has both the largest latency and least bandwidth of all GPU memories. For efficient kernels it is important to minimize memory read and writes from and to HBM.
Specialized Compute Units. Modern GPUs have specialized compute units – called tensor cores – that accelerate matrix multiplications on GPUs. Tensor cores have most of the GPU compute and are accessed at the warp or block level. For efficient kernels it is important to maximize tensor core utilization.
@@ 3.2 More Efficient Kernels via Two Level Sequence Parallelism
TiledFlashLinearAttention (TFLA) enables fast kernels and a trade-off between memory consumption and computational efficiency by introducing two levels of sequence parallelism (see Figure 6). The first level is the parallelization over the chunks of the sequence, which requires to compute and materialize intermediate states $\bm{C}_{k}$ in GPU HBM. For this we use a recurrent kernel similar to previous work (Yang et al., 2024b). The second level is the parallelization within each chunk, which is achieved by tiling the intra chunk attention matrix along the chunk dimension. This second level of parallelism enables large chunk sizes and hence reduces the memory consumption for the intermediate states as we have to store and load $N_{c}=\lceil T/L\rceil$ intermediate states in HBM on each kernel call, where $T$ is the sequence length and $L$ is the chunk size. In addition to the two levels of sequence parallelism and the naive parallelization over the batch and head dimensions, TFLA also parallelizes over the embedding dimension. This enables arbitrary large head dimensions and results in a massive parallelization over five dimensions, which is crucial for achieving high performance on modern GPUs.
@@
In simplified form, the intra-chunk parallel forward pass of the mLSTM (and other linear RNNs) for a chunk $k$ can be written as three matrix multiplications, which we fuse into a single kernel:
@@
In Appendix A.3, we show that TFLA can be applied to any linear RNN that either follows or can be reformulated into this form. In order to parallelize the computation in (20), we introduce the block sizes $B_{Lhq}$, $B_{Lkv}$, $B_{dqk}$ and $B_{dhv}$ for the attention matrix, query, key, value and hidden state dimensions $L_{hq}$, $L_{kv}$, $d_{qk}$ and $d_{hv}$, along which we either parallelize or accumulate by using a loop inside the kernel.
In Figure 3, we show our TFLA tiling strategy for the forward pass $\bm{\mathrm{H}}^{(k)}$ kernel. We parallelize across the outer sequence dimension $L_{hq}$ with $N_{Lhq}=L_{hq}/B_{Lhq}$ programs, and across the outer embedding dimension $d_{hv}$ with $N_{dhv}=d_{hv}/B_{dhv}$ programs. We loop over the inner dimensions $L_{kv}$ and $d_{qk}$, which are tiled by the block sizes $B_{Lkv}$ and $B_{dqk}$ respectively.
@@
For the mLSTM we cannot simply accumulate the results of the matrix multiplications $\bm{\mathrm{H}}^{(k)}_{\text{intra}}$ along the $L_{kv}$ dimension and $\bm{\mathrm{H}}^{(k)}_{\text{inter}}$ due to the stabilization of the exponential input gate with the max state $m_{t}$. The max state tracks the maximum of the forget and input gates over time and is used to stabilize the exponential input gate similar to the safe softmax computation (Milakov & Gimelshein, 2018). Since we compute the hidden state output $\bm{\mathrm{H}}^{(k)}$ in blocks along the chunk size (i.e. time) dimension $L_{kv}$, we need to rescale during accumulation of the block results for $\bm{\mathrm{H}}^{(k)}_{\text{intra}}$ and the overall results into $\bm{\mathrm{H}}^{(k)}$ in the same way as FlashAttention (Dao, 2024). We provide details on the rescaling in Section B.2. For the backward pass there is no rescaling necessary as we store the max states in the forward pass and reuse them in the backward pass. The pseudocode for the forward pass of TFLA for the mLSTM is listed in Algorithm 1.
@@
$\delta\bm{Q}^{(k)}$ P L P L $\left(\frac{d_{qk}}{B_{dqk}},\frac{L_{hq}}{B_{Lhq}},\dots\right)$
$\delta\bm{K}^{(k)}$ L P P L $\left(\frac{d_{qk}}{B_{dqk}},\frac{L_{kv}}{B_{Lkv}},\dots\right)$
$\delta\bm{V}^{(k)}$ L P L P $\left(\frac{d_{hv}}{B_{dhv}},\frac{L_{kv}}{B_{Lkv}},\dots\right)$
@@ 4.2 Normalization of mLSTM and Linear RNNs
Motivated by the performance of mLSTMsig, we seek to understand the differences between mLSTMsig and mLSTMexp empirically. To approach this, we draw inspiration from the concept of frequency response and transfer function analysis for control systems design, where typically the amplitude ratio or gain of output and input signals for different frequencies is considered (Ogata, 2010, Ch. 7). In our case, we analyze the transfer behavior of mLSTMsig and mLSTMexp for random inputs $\bm{q}_{t},\bm{k}_{t}$ and $\bm{v}_{t}$ and different input gate and forget gate preactivations $\tilde{\mathrm{i}}_{t}$ and $\tilde{\mathrm{f}}_{t}$.
@@
Interestingly, almost all other gated linear RNN variants also place a normalization layer after the RNN cell (Dao & Gu, 2024; Sun et al., 2023; Qin et al., 2024b; Yang et al., 2024b). Often this is justified with improved training stability, but a more thorough discussion is missing (Lieber et al., 2024). Qin et al. (2022) analyze the effect of the norm layer after a non-gated, kernel-based linear attention layer (Katharopoulos et al., 2020) and show that this effectively prevents unbounded gradients. We also confirm that the norm layer has a significant impact on training stability and the gradient norm during training. In Section 5.1 we show that initializing the input gate bias at larger negative values, as suggested by our transfer behavior analysis in Figure 4, prevents large gradient norm variance and spikes during training.
@@
We hypothesize that at this point the normalization layer does not only have a stabilizing effect by controlling the magnitude of the layer activations through rescaling, but also actively participates in the information routing or gating mechanism of the linear RNN. For example, if the squared norm of $\bm{C}_{t}^{\top}\,\bm{q}$, which is controlled by input and forget gates through $\bm{C}_{t}^{\top}$, is smaller than the epsilon, the denominator in the $\mathrm{NORM}(\bm{x})$ layer is dominated by $\epsilon$ and the output moves towards zero (indicated by the purple area in Fig. 4). Hence, by moving through the x-y plane in Fig. 4, the gates could learn to suppress or amplify any input in the sequence. In Section D.2 we show additional experiments on the effect of varying the normalization layer epsilons and different modifications of the normalizers for the mLSTM.